OpenStudy (lindsey11x16):

If I am given the functions: tanx = root3/3 and cosx = -root3/2 how do i find sinx ?

OpenStudy (anonymous):

\[\tan =\frac{ \sin }{ \cos } (=) \sin x=\cos x \tan x (=) \sin x = -\frac{ \sqrt{3} }{ 2 } \times \frac{\sqrt{3} }{ 3 } \] \[\sin x = -\frac{ \sqrt{3}^{2} }{ 6 } (=) sinx =- \frac{ 3 }{ 6 } (=) \sin x =-\frac{ 1 }{ 2 }\]

OpenStudy (lindsey11x16):

wow that made complete sense if i i wanted to look for cosecant, it would be -2/1 ? is this correct?

OpenStudy (anonymous):

give me a sec, it been a while since i worked with cosec

OpenStudy (lindsey11x16):

sure!

OpenStudy (anonymous):

yes that is correct since \[cosec = \frac{ 1 }{ \sin x }\]

OpenStudy (lindsey11x16):

Thank you!